# LoVA: Long-form Video-to-Audio Generation

### v1

@@ I Introduction
Video-to-Audio (V2A) generation, which aims to create synchronized and realistic sound effects for silent videos, finds widespread use in video editing, sound effect creation, and autonomous content enhancement [1]. However, current V2A methods predominantly focus on generating fixed-length short audios, typically less than 10 seconds on VGGSound [2] or Audioset [3] benchmarks. These methods generate fixed-length audios through autoregressive approaches truncated to a maximum length [4, 5, 6], or fixed-length noise denoising by UNet-based diffusions [7, 8, 9]. Despite their success in generating fixed-length short audios, the challenge of creating audio for variable-length, long-form videos exceeding 10 seconds in real-world scenarios remains unexplored. Our work aims to address this short-to-long duration gap in the V2A domain.
Figure 1: Long-form V2A example. Current (8s/10s) UNet-based diffusion V2A models (DiffFoley, TiVA, FoleyCrafter) exhibit inconsistency when generating long-form (30s) audio, as indicated by clear mel-spectrogram boundaries and structural variances. In contrast, our LoVA produces consistent results akin to the ground truth.
@@
To address the aforementioned challenge, we introduce LoVA, a Long-form Video-to-Audio generation model adept well at handling long-duration problem. As depicted in Figure 2(a), the expected long-form V2A model should possess the capabilities of: (1) maintaining the variable-length audio as a sequence of lossless frames to ensure quality; (2) modeling the full sequence interactions among frames, rather than the localized interactions learned by convolutional UNets, to ensure feasibility and consistency when extending to long sequences; (3) generating multiple frames in parallel for efficiency. Diffusion Transformer (DiT)[16] treats latent data as token sequences in the diffusion process, aligning well with the sequential nature of audio. It also demonstrated promising results in generating high-quality images[16], videos [17], and audios [18, 19] in an efficient parallel sequence generation manner. Thus, we introduce DiT into the V2A domain and model the denoising process on noisy latent audio frames, termed as LoVA. For long-form V2A problems, LoVA simply extracts extended video features and prepares correspondingly longer sequences of noisy audio frames for denoised lengthy audio generation, akin to a consistent frog croak over a 30s video as depicted in Figure 1.
@@
Furthermore, acknowledging the absence research in the long-form V2A domain, we have established a long-form V2A evaluation based on a variable-length long video dataset UnAV100 [20], as an addition to the current standard short-form evaluation. We conducted extensive experiments on this long-form evaluation to validate the performance of autoregressive, UNet-based diffusion, and DiT-based diffusion (our LoVA) methods, as well as their duration-extending characteristics in long-form V2A.
@@ II-B Architecture of LoVA
where $N$ is the frame number and $h_{C}$ is the CLIP hidden size. To further assist the model in learning the relationship between the audio sequence and video frames when extending to long-form V2A generation, we add a learnable positional embedding layer $PE_{c}$ to the video condition $c$:
@@ III-D Metrics
Besides, we introduce the number of inferences per audio (Num.Infer.) as a indicator of potential inconsistency. We also randomly select 40 videos from UnAV100 test set for human evaluation. Evaluators are asked to give a 5-level Likert scale on 4 aspects:Sound Quality (SoundQua.), Semantic Relevance (Sem.Rel.), Consistency (Cons.) and Overall quality (Overall).
@@ IV-B Comparison between Different Diffusion Denoiser
We conduct ablation studies to validate the effectiveness of DiT modules when performing long-form V2A generation. Being similar with LoVA, some previous UNet-based diffusion models, like FoleyCrafter [8], can also generate long-form audio by modifying the shape of the latent space. On the UnAV100 benchmark, we cut the video into clips with different durations and generate audio for each clip individually.Then we obtain the long-form audio by concatenating all generated audio clips. FoleyCrafter is adapted to different clip durations by resizing the latent space.
Figure 3: Comparison of long-form audio generation ability between UNet and DiT structure. We choose FAD and KL to represen generated audios’ quality. The experimennt is carried on UnAV100 test dataset. Different splitting duration means different sub-video durations, as well as different generated sequence lengths per inference.
As shown in Fig 3, as the clips duration increases, the FAD and KL metrics degenerate. The best scores are achieved when the splitting duration is set to 10 seconds, which aligns with the training data on which FoleyCrafter were trained. However, for DiT-based LoVA (w/o tuning), metrics do not show obvious degenerative phenomenon as the duration increases. It should be stressed that FoleyCrafter and LoVA are all trained on 10-second data only, but performs differently on long-form audio generation. This difference highlights a critical limitation of the UNet structure when it extends to long-form audio, while proves DiT’s effectiveness to handle long audio sequence.

### v2

@@ I Introduction
Video-to-Audio (V2A) generation, which aims to create synchronized and realistic sound effects for silent videos, finds widespread use in video editing, sound effect creation, and autonomous content enhancement [1]. However, current V2A methods predominantly focus on generating fixed-length short audios, typically less than 10 seconds on VGGSound [2] or Audioset [3] benchmarks. These methods generate fixed-length audios through autoregressive approaches truncated to a maximum length [4, 5, 6], or through UNet-based diffusions to denoise fixed-length noise [7, 8, 9]. Despite their success in generating fixed-length short audios, the challenge of creating audio for variable-length, long-form videos exceeding 10 seconds in real-world scenarios remains unexplored. Our work aims to address this short-to-long duration gap in the V2A domain.
Figure 1: Long-form V2A example. Current (8s/10s) UNet-based diffusion V2A models (DiffFoley, TiVA, FoleyCrafter) exhibit inconsistency when generating long-form (30s) audio, as indicated by clear mel-spectrogram boundaries and structural variances. In contrast, our LoVA produces consistent results similar to the ground truth.
@@
To address the aforementioned challenge, we introduce LoVA, a Long-form Video-to-Audio generation model that is designed to handle long-duration problem. As depicted in Figure 2(a), the expected long-form V2A model should possess the capabilities of: (1) maintaining the variable-length audio as a sequence of lossless frames to ensure quality; (2) modeling the full sequence interactions among frames, rather than the localized interactions learned by convolutional UNets, to ensure feasibility and consistency when extending to long sequences; (3) generating multiple frames in parallel for efficiency. Diffusion Transformer (DiT)[16] treats latent data as token sequences in the diffusion process, aligning well with the sequential nature of audio. It also demonstrated promising results in generating high-quality images[16], videos [17], and audios [18] in an efficient parallel sequence generation manner. Thus, we introduce DiT into the V2A domain and model the denoising process on noisy latent audio frames, termed as LoVA. For long-form V2A problems, LoVA simply extracts extended video features and prepares correspondingly longer sequences of noisy audio frames for denoised lengthy audio generation, akin to a consistent frog croak over a 30s video as depicted in Figure 1.
@@
Furthermore, acknowledging the absence of research in the long-form V2A domain, we have established a long-form V2A evaluation based on a variable-length long video dataset UnAV100 [19], as an addition to the current standard short-form evaluation. We conducted extensive experiments on this long-form evaluation to validate the performance of autoregressive, UNet-based diffusion, and DiT-based diffusion (our LoVA) methods, as well as their duration-extending characteristics in long-form V2A.
@@ II-B Architecture of LoVA
where $N$ is the number of frames and $h_{C}$ is the CLIP hidden size. To further assist the model in learning the relationship between the audio sequence and video frames when extending to long-form V2A generation, we add a learnable positional embedding layer $PE_{c}$ to the video condition $c$:
@@ III-D Metrics
Besides, we introduce the number of inferences per audio (Num.Infer.) as an indicator of potential inconsistency. We also randomly select 40 videos from UnAV100 test set for human evaluation. Evaluators are asked to give a 5-level Likert scale on 4 aspects:Sound Quality (SoundQua.), Semantic Relevance (Sem.Rel.), Consistency (Cons.) and Overall quality (Overall).
@@ IV-B Comparison between Different Diffusion Denoiser
We conduct ablation studies to validate the effectiveness of DiT modules when performing long-form V2A generation. Being similar to LoVA, some UNet-based diffusion models, like FoleyCrafter [8], can also generate long-form audio by expanding the shape of latent space. On the UnAV100 benchmark, we split the video into sub-videos with different splitting durations and generate sub-audio for each sub-video individually. Then we obtain the long-form audio by concatenating all generated sub-audios. FoleyCrafter is adapted to different splitting durations by resizing the shape of latent space.
Figure 3: Comparison of long-form audio generation ability between UNet and DiT structure. We choose FAD and KL to represent generated audios’ quality. The experimennt is carried on UnAV100 test dataset. Different splitting durations mean different sub-videos’ and generated sub-audios’ durations per inference.
As shown in Fig 3, as the splitting duration increases, the FAD and KL metrics degenerate. The best scores are achieved when the splitting duration is 10 seconds, aligning with the training data’s duration of FoleyCrafter. However, for DiT-based LoVA (w/o tuning), metrics do not exhibit obvious degeneration phenomenon as splitting duration increases. Notably, both FoleyCrafter and LoVA are trained on 10-second data only, yet perform differently on long-form audio generation. This difference highlights a critical limitation of the UNet structure when it extends to long-form audio, while proves DiT’s effectiveness to handle long audio sequence.
@@ V Conclusion and Future Work
Acknowledgments
This work is supported by the National Natural Science Foundation of China (No.62276268) and ZHI-TECH GROUP.